OpenStudy (toxicsugar22):

According to the following reaction, how many moles of zinc sulfate will be formed upon the complete reaction of 0.941 moles sulfuric acid with excess zinc hydroxide? H2SO4 (aq) + Zn(OH)2 (s) = ZnSO4 (aq) + 2 H2O (l)

OpenStudy (anonymous):

ok 1 H2SO4 = 1 ZNSO4 0.941 H2S04 = X mole ZNS04 THEREFORE ,mole of H2S04= 0.941 MOLES

OpenStudy (anonymous):

0.941

OpenStudy (anonymous):

Absolutely affirmative

OpenStudy (anonymous):

4*0.215=Ans

OpenStudy (anonymous):

Look at the equation look for the ratio of P(4):PCl(3) which is 1:4 that means 1 P4=4 PCl3 0.251 P4= X PCl3 which means X PCl3= 4*0.251 which means moles of PCl3= 4*0.251,which is i am sure your answer you know why bc i have confidence in you

OpenStudy (anonymous):

sorry you are right

OpenStudy (anonymous):

i used 0.251 instead of 0.215

OpenStudy (anonymous):

0.86

OpenStudy (toxicsugar22):

According to the following reaction, how many moles of ammonium nitrate are necessary to form 0.177 moles water? NH4NO3 (aq) N2O (g) + 2 H2O (l)

OpenStudy (anonymous):

you want me to do this too

OpenStudy (anonymous):

1 NH4NO3 = 2 H2O X NH4NO3 = 0.177 H2O therefore moles of NH4NO3=(0.177/2) moles

OpenStudy (anonymous):

88.5 * 10^-3

OpenStudy (anonymous):

have you seen it

OpenStudy (toxicsugar22):

.0885 is the answer

OpenStudy (toxicsugar22):

.0885

OpenStudy (anonymous):

88.5 * 10^-3 not to the power 3

OpenStudy (toxicsugar22):

sowhat is the answer

OpenStudy (toxicsugar22):

is it .0885

OpenStudy (anonymous):

88.5 * 10^-3=0.885 are the same thing

OpenStudy (toxicsugar22):

thanks